Ticket Number: HEA-NOC/20040126-5 Ticket Status: CLOSE
Ticket Type: scheduled Resolver: HEA-NOC
Ticket Opened: 20040126 12:27 UTC Problem Start: 20040128 08:45 UTC
Ticket Closed: 20040130 14:18 UTC Problem End: 20040128 09:00 UTC
The primary UCD access circuit will be resized to 190 Mbps. Also during
this window, the fibre patching for the circuit in the HEAnet Brooklawn
House POP will be rerouted.
All traffic to/from UCD will reroute over a secondary link during the work,
with impeded performance.
20040126 12:24 BB Scheduling work.
20040128 10:27 BB The circuit has been upgraded. There was a problem with
the router on the HEAnet side, which may or may not be related to this. It
will be tracked in http://www.hea.net/tickets/20040128-1. Otherwise,
monitoring circuit for 24 hours.
20040130 14:17 BB Circuit has been stable for 24 hours. Upgrade to 190
Time to Fix:
This ticket can be monitored at http://www.hea.net/tickets/20040126-5